Circular No. 7786

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002A IN UGC 3804 Y. L. Qiu and J. Y. Hu, Beijing Astronomical Observatory (BAO),

report the discovery by the BAO Supernova Survey of an apparent supernova in unfiltered images taken with the BAO 0.6-m telescope on Jan. 1.79 (mag ∼ 17.4), 8.61 (mag ∼ 16.8), and 9.56 UT (mag ∼ 16.8). SN 2002A is located at α = 7^{h}22^{m}36^{s}.14, δ = +71^{o}35′41′′.5 (equinox 2000.0), which is 6′′.2 east and 14′′.7 south of the nucleus of UGC 3804. An image taken with the same telescope on 2000 Dec. 27 showed no star at this position (limiting mag ∼ 18.7).

PECULIAR VARIABLE IN MONOCEROS M. Della Valle, Osservatorio Astrofisico di Arcetri, Florence; and T.

Iijima, Osservatorio Astronomico di Padova, write: “Preliminary analysis of spectra (range 400–680 nm, resolution ∼ 0.5 nm) of the variable star reported on IAUC 7785, obtained on Jan. 8 with the 1.22-m telescope (+ Boller & Chivens spectrograph), shows a heavily reddened continuum characterized by a lack of the typical signatures exhibited by classical novae at maximum light or during their early decline. No absorption band of TiO or C_{2} is seen, while we detect faint Hα and Na I emission lines flanked by P-Cyg profiles. In addition, we note the presence of three narrow emission lines at 650.2, 614.5, and 585.9 nm, also flanked by P-Cyg profiles, tentatively identified as Ba II. The expansion velocities derived from the minima of the P-Cyg profiles range between 380 and 500 km/s. All of this may suggest this object not to be a normal classical nova, but rather a peculiar ‘slow’ nova or a post-asymptotic-giant-branch star in flare-up. Knowledge of the amplitude of the outburst could improve significantly the classification.”

T. Kato, Kyoto University, forwards the following position for the variable measured by H. Yamaoka (Kyushu University) from a V -band CCD image taken on Jan. 9.496 UT by S. Kiyota (Tsukuba, Ibaraki, Japan; 0.25-m reflector): α = 7^{h}04^{m}04^{s}.801, δ = −3^{o}50′50′′.77 (equinox 2000.0; 29 Tycho-2 stars, uncertainty ∼ 0′′.3). They suggest that this object is identical to IRAS 07015–0346, or the 2MASS source having position end figures are 04^{s}.82, 50′′.5 (and magnitudes J = 13.905, H = 13.480, K = 13.347), noting that the infrared and V − J colors exclude the possibility of a long-period variable (and that the IRAS color is very unusual for a stellar object). Kiyota provides the following additional magnitudes: Jan. 9.497, I_{c} = 8.28; 9.498, V = 10.10, B = 12.16.

2002 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green